OpenStudy (anonymous):

Connie has saved up $15 to purchase a new CD from the local store. The sales tax in her county is 5% of the sticker price. Write an equation and solve it to determine the value of the highest priced CD Connie can purchase with her $15, including the sales tax. Round your answer to the nearest penny. x + 0.05x = 15; x = $14.29 x - 0.05x = 15; x = $15.79 0.5x = 15; x = $30 5x = 15; x = $3.00

OpenStudy (anonymous):

help!!!

OpenStudy (anonymous):

I remember solving this question

OpenStudy (anonymous):

Yeah XD i solved it with you

OpenStudy (anonymous):

It should be the second one

OpenStudy (anonymous):

i dont beleive i asked this

OpenStudy (anonymous):

it wasnt the second one apparently

OpenStudy (ranga):

Let x be the price of a CD. The tax on that will be 5% of x or x * 0.05 Total cost = cost of CD + tax = x + 0.05x The equation is x + 0.05x = 15 Solve for x: x + 0.05x = 15 1.05x = 15 x = 15/1.05 = $14.29